### v2

# Spectral Bottleneck in Sinusoidal Representation Networks: Noise is All You Need

Abstract
This work identifies and attempts to address a fundamental limitation of implicit neural representations with sinusoidal activation. The fitting error of SIRENs is highly sensitive to the target frequency content and to the choice of initialization. In extreme cases, this sensitivity leads to a spectral bottleneck that can result in a zero-valued output. This phenomenon is characterized by analyzing the evolution of activation spectra and the empirical neural tangent kernel (NTK) during the training process. An unfavorable distribution of energy across frequency modes was noted to give rise to this failure mode. Furthermore, the effect of Gaussian perturbations applied to the baseline uniformly initialized weights is examined, showing how these perturbations influence activation spectra and the NTK eigenbasis of SIREN. Overall, initialization emerges as a central factor governing the evolution of SIRENs, indicating the need for adaptive, target-aware strategies as the target length increases and fine-scale detail becomes essential. The proposed weight initialization scheme (WINNER) represents a simple ad hoc step in this direction and demonstrates that fitting accuracy can be significantly improved by modifying the spectral profile of network activations through a target-aware initialization. The approach achieves state-of-the-art performance on audio fitting tasks and yields notable improvements in image fitting tasks.
I Introduction
Figure 1: Spectral bottleneck issue shown using SIREN and overview of the weight perturbation scheme. (a) We attempt to fit eight discretely sampled broadband 1D signals (S1–S8) with decreasing low-frequency content. As shown in the right panel, the PSNR of SIREN sitzmann2020implicit progressively decreases from S1 to S4, eventually encountering a spectral bottleneck after S4. SIREN fails to capture nearly all the frequencies of signals S4–S8, even though its frequency support should in principle allow it to represent a portion of the spectrum. In contrast, SIREN2 initialized with WINNER maintains higher PSNR across all signals. (b) Schematic of a feedforward neural network with periodic activations, illustrating the statistical distributions of layer-1 outputs $X_{ij}$, weight matrix $W_{jk}$, and noise matrix $\eta_{jk}$. The effect of Gaussian noise ($\eta_{jk}$) on the spectrum of layer 2 pre-activations is shown: WINNER enhances the receptivity of high-frequencies.
Implicit neural representations (INRs) of natural and synthetic signals have broad applications across various domains. They allow neural networks to represent coordinate-based discrete data such as images, videos, audio, 3D shapes, and scientific datasets as continuous functions. This allows seamless integration of multimedia, medical, and scientific data into machine learning pipelines for tasks such as denoising, classification, inpainting, and latent representation mildenhall2021nerf; oechsle2019texture; niemeyer2020differentiable; szatkowski2022hypersound. The key advantages of INRs over discretely sampled data are continuous input parameterization (no grid), fully differentiable networks with accessible gradients, and compressed data representation sitzmann2020implicit; tancik2020fourier. Using the readily available spatio-temporal gradients obtained through automatic differentiation, INRs can also be employed to solve forward and inverse problems governed by differential equations in a mesh-free setting raissi2019physics; song2022versatile.
Although multi-layer perceptrons (MLPs) are universal function approximators, training them to fit natural signals such as images, videos, or audio is often challenging. Analytical and empirical results are presented in previous studies rahaman2019spectral; xu2019frequency showing that ReLU based deep neural networks are prone to ‘spectral bias’ and are ill-conditioned for low-dimensional coordinate-based training tasks. To overcome spectral bias and the associated lazy training, coordinate inputs are often mapped to positional encodings mildenhall2021nerf or random Fourier features rahimi2007random; tancik2020fourier, which as shown in yuce2022structured increase the ‘expressive power’ of INRs. Alternatively, Sitzmann et al. sitzmann2020implicit have proposed sinusoidal representation networks (SIRENs) using periodic activation function $\phi(x)=\sin(\omega_{0}x)$,
$f^{\text{SIREN}}(\mathbf{x};\theta)$ $=\mathbf{W}^{(L)}\mathbf{h}^{(L-1)}+\mathbf{b}^{(L)},$ (1)
$\mathbf{h}^{(l)}$ $=\phi^{\text{sin}}\!\left(\mathbf{W}^{(l)}\mathbf{h}^{(l-1)}+\mathbf{b}^{(l)}\right),\quad\mathbf{h}^{(0)}=\mathbf{x}.$
where $\mathbf{x}\in\mathbb{R}^{d}$ is the input, $\theta=\{\mathbf{W}^{(l)},\mathbf{b}^{(l)}\}_{l=1}^{L}$ are the network parameters, L1L-1 hidden layers, and $\omega_{0}$ is the activation periodicity specified as hyperparameter. They also propose a principled initialization scheme for the weights $\mathbf{W}$, to ensure network’s pre- and post-activations at initialization remain narrowly bounded. Their approach faithfully captures both the discrete data and its gradients with high-fidelity, enabling applications in computer vision and the solution of scientific differential equations. Various other variants related to SIREN were also proposed, such as ziyin2020neural; liu2013fourier; silvescu1999fourier.
While SIRENs with the uniform weight initialization scheme sitzmann2020implicit can mitigate spectral bias and effectively fit 2D images, videos, and 3D geometries, they often perform poorly on low-dimensional signals such as audio when low-frequency components contribute minimally to the target. As pointed out in yuce2022structured, the capacity of network to represent a target signal’s frequency components does not guarantee efficient learning of that signal. The reconstruction accuracy is rather closely tied to the spectral content of the signal and the inductive bias of network initialization; for instance, under fixed initialization, SIRENs struggle with signals dominated by very high or very low frequencies. Prior work mehta2021modulated; liu2024finer shows that SIRENs exhibit overfitting as signal length and high-frequency content increase. An easy way to get around this problem is to increase the input dimensionality by mapping the input coordinates to a random Fourier feature space rahimi2007random; tancik2020fourier; mildenhall2021nerf. However, positional embeddings lead to a quadratic increase in the parameter count with respect to the embedding dimension, and consequently with the hidden-layer width. In this work, we examine the fitting challenges of SIREN on increasingly high-frequency targets and characterize the limitations of its traditional weight initialization sitzmann2020implicit, particularly the spectral bottleneck phenomenon, where SIREN fails to recover all frequency modes and collapses to a zero-valued function. We then propose a noisy weight initialization scheme WINNER that aims to mitigate the inductive bias at initialization. Figure 1 illustrates the spectral bottleneck issue of the standard SIREN sitzmann2020implicit and the proposed weight initialization scheme which is being employed in SIREN2. Representation accuracy of SIREN is compared with the proposed method on several synthetically generated audio signals, each containing 150,000 samples. Although, the spectral bottleneck phenomenon in Fig. 1 is demonstrated through synthetic signals, such a phenomenon was noted to also observe in fitting natural signals, examples include tap.wav and relay.wav audio clips which can be accessed through the project’s repository (link in first page). All experiments in Figure 1 use a SIREN architecture with four hidden layers with 222 features in each layer.
The key contributions of this work are:
  1. 1.
    Show that SIRENs can suffer from a spectral bottleneck, and analyze their training dynamics in Fourier space to understand how this phenomenon unfolds during training.
  2. 2.
    A new target-aware weight perturbation scheme WINNER, that adds noise into the baseline uniformly distributed weights of SIRENs to broaden their frequency support according to the target and thereby avoid spectral bottleneck.
  3. 3.
    The influence of the proposed noise addition scheme on the spectral distribution of pre-/post-activations and the eigenbasis of the empirical Neural Tangent Kernel (NTK) at initialization is analyzed.
II Related work
Activation functions for INRs. Sinusoidal activation functions have been considered in neural networks since 1999 sopena1999neural, but such models were often regarded as difficult to train parascandolo2016taming. Sitzmann et al. sitzmann2020implicit addressed this by introducing SIREN, which uses a tailored initialization scheme that enhances stable training of sinusoidal INRs. Later methods designed alternative periodic or adaptive activations jagtap2022deep to extend frequency support, such as FINER liu2024finer, FINER++ zhu2024finerplusplus, and HOSC serrano2024hosc, while others explored Gaussian or kernel-inspired nonlinearities ramasinghe2022beyond; chng2022gaussian. DINER zhu2024disorder further demonstrated that near-lossless signal representation can be achieved by augmenting a hash-table to a traditional INR backbone albeit with an additional memory overhead to store the hash data. Additional activation designs include Snake, a periodic trainable activation for learning periodic targets liu2020neural; WIRE, which uses Gabor wavelet activations to capture frequency and spatial locality saragadam2023wire; a sampling-theoretic case for using $\mathrm{sinc}$ activations in INRs saratchandran2024sampling; and STAF, a family of trainable sinusoidal activations morsali2025staf.
Weight/bias initialization. Recent works like FINER liu2024finer and FINER++ zhu2024finerplusplus explore bias initialization strategies to reduce eigenvalue decay in the empirical NTK, thus increasing effective frequency support. In geometry-focused INRs, SAL initializes network weights so that the signed distance function (SDF) starts as a 3D sphere atzmon2020sal; DiGS later introduced a multifrequency geometric initialization (MFGI) that extends this idea to sine activations and preserves high-frequency content ben2022digs; and in sinusoidal INRs, SIREN showed that without a tailored scheme, sine activations fail to reconstruct signals, motivating an initialization that stabilizes optimization sitzmann2020implicit. The works of Tang et al. tang2025structured and Varre et al. varre2023spectral (for linear networks) demonstrate how initialization can govern parameter optimization. Complementary strategies have been developed for sinusoidal and general INRs: VI3NR derives an activation-agnostic, variance-preserving rule for INR regression hewa2025vi3nr; TUNER introduces frequency-sampling input initialization together with amplitude-bounded hidden-layer weights to control the model bandlimit during training novello2025tuning; and Yeom et al. yeom2024fast propose an initialization scaling for sinusoidal representation networks to speed up optimization and alleviate spectral bias. Prior designs are often specialized (geometry for SAL/DiGS and sine activations for SIREN) and typically do not target the joint distribution of pre- and post-activations across layers. In this work, we introduce a target aware weight initialization scheme that takes into account the spectral characteristics of the target as well as preserves the layer distributions for stability and accuracy across INR tasks.
Audio INRs. A variety of models have been proposed to enhance the ability of neural networks to represent audio signals choudhury2024nerva; kim2023regression; kim2023generalizable; li2024asmr. Techniques such as hypernetworks, positional encoding, and auxiliary networks have also been explored to improve reconstruction fidelity and reduce reconstruction noise marszalek2025hypernetwork; lanzendorfer2023siamese. Audio-specific works include Siamese SIREN lanzendorfer2023siamese for compression, HyperSound szatkowski2022hypersound for INR generation via meta-learning, and INRAS su2022inras for spatial audio modeling. Neural audio representations continue to find applications in classification, speech synthesis, sound event detection, encoding, and embedding sharan2021benchmarking; tessarini2022audio; hershey2017cnn; donahue2018adversarial; shen2018natural; aironi2021graph.
In addition, analytical and empirical analysis of Basri et al. basri2020frequency; ronen2019convergence indicate that for 1D signals, the convergence rate under gradient descent scales inversely with the square of the target signal’s frequency (i.e., $1/k^{2}$), and this frequency-dependent slowdown grows exponentially with the input dimension bietti2019inductive; cao2019towards. This frequency-dependent limitation has motivated several INR-specific methods. For example, FR-INR reparameterizes MLP weights as linear combinations of fixed Fourier bases to mitigate low-frequency bias, while IGA estimates a gradient transform from a sampled empirical NTK and applies it during updates to accelerate learning of high-frequency components shi2024frinr; shi2024iga.
III Understanding ‘Spectral Bottleneck’
III-A Challenges of Fitting 1D Signals with SIREN
To maintain uniform feature scales and prevent exploding gradient issues, it is standard practice to normalize network inputs during training. However, for 1D data with high sampling rates, such as audio signals, input normalization introduces a mismatch between the frequency content of the signal and the frequency range effectively supported by the network. For instance, the input normalization $\mathbf{x}\sim\mathcal{U}(-1,1)$, scales the maximum frequency of a discretely sampled signal in proportion to its sampling rate.
As for SIREN sitzmann2020implicit, increasing the activation periodicity $\omega_{0}$ (Eqn 1) may seem like a potential solution to mitigate the frequency bias induced by input scaling. However, since the pre-activation values of a SIREN predominantly lie within the range of [3,3][-3,3] due to its weight initialization scheme, a high $\omega_{0}$ can render the activation function insensitive and alter the activation distribution. To address this, Sitzmann et al. sitzmann2020implicit scaled the inputs by a factor of 100, sampling them from a wider range $\mathbf{x}\sim\mathcal{U}(-100,100)$, which effectively increases the power spectral density of the pre-activations by approximately the same factor (see Proposition 1). While this strategy is effective for low-frequency-dominant signals, SIREN with input scaling can still fail on high-frequency-dominant signals. An example illustrating this failure is discussed in Fig. 2.
Figure 2: An example case where the standard weight initialization scheme of SIREN fails to reconstruct an audio clip. (a) Spectrograms of ground truth (tetris.wav) (top), SIREN reconstruction (middle), and the error map (bottom) computed using first layer $\omega_{0}=30000$ and hidden layer width 128. (b,c) PSNR histories of SIREN for different input scalings and network sizes. A five-layer MLP was used with a learning rate scheduler reducing the rate by 2% every 20 epochs from an initial value of $10^{-4}$.
Proposition 1.
If $\mathbf{w}_{a}\sim\mathcal{U}(-a,a)^{d}$, then $z_{a}(\mathbf{x})=\mathbf{w}_{a}^{\top}\mathbf{x}$ can be written as $a\,z_{1}(\mathbf{x})$ with $\mathbf{w}_{1}\sim\mathcal{U}(-1,1)^{d}$, hence its Fourier transform scales by $a$ and the power spectral density satisfies $S_{a}(k)=a^{2}S_{1}(k)$.
An example case when SIREN fails to fit a high-frequency signal: We consider the reconstruction of a high-frequency audio signal $f:\mathbb{R}^{2}\to\mathbb{R}^{d}$, using the example tetris.wav available in the linked GitHub repository. Experiments conducted using SIREN in Fig. 2(a,b) show that, although input scaling improves PSNR value, significant errors remain in the reconstructed signal, as shown in the spectrogram error. A maximum PSNR value of just $\sim$23.5dB was achieved using a scaling factor of $3\times 10^{4}$. We also experimented by increasing network size (Fig. 2b), changing scheduler parameters, and changing the frequency parameter $\omega_{0}$, none of which resolve this issue. This reveals that SIREN, with its default weight initialization, struggles to fit such high-frequency dominant signals due to the spectral bottleneck phenomenon. Other 1D examples are shown in Fig. 1b, where SIREN exhibits a spectral bottleneck for signals S4–S8, all lacking low-frequency components.
Remark. Repeating the experiment in Fig. 2 with alternative architectures, including WIRE saragadam2023wire, FINER liu2024finer, HOSC serrano2024hosc, and Gauss ramasinghe2022beyond, all yield the same outcome: convergence to a fixed PSNR of 13.4 with a zero-valued output. This pathology is therefore not just confined to SIRENs, but is shared across other related deep neural networks. We further observed that mapping the inputs to random Fourier features rahimi2007random; tancik2020fourier or adopting a broader bias initialization, as proposed in zhu2024finerplusplus, mitigates this failure mode. A formal characterization of this behavior is presented in Sec. III-B.
III-B Learning Dynamics in Fourier Space
To explore the optimization trajectory that leads to the spectral bottleneck associated with SIRENs, as observed in Fig. 2, we examine the network’s learning dynamics in Fourier space through the lens of its empirical Neural Tangent Kernel (NTK) rahaman2019spectral; wang2022and; basri2020frequency; tancik2020fourier. Before analyzing the training dynamics, we first define the NTK eigenbasis and discuss its interpretation.
Jacot et al. jacot2018neural showed that in the infinite-width limit, fully connected networks trained with infinitesimal learning rates follow linear training dynamics governed by the NTK. The NTK is a Gram matrix defined as,
$\Theta(\mathbf{x},\mathbf{x}^{\prime})=\nabla_{\theta}\Phi(\mathbf{x};\theta)\cdot\nabla_{\theta}\Phi(\mathbf{x}^{\prime};\theta),$ (2)
where $\Phi(\mathbf{x};\theta)$ is the network output and $\theta$ its parameters. In the infinite-width regime, $\Theta$ remains constant during training, allowing for a closed-form linear model of training dynamics lee2019wide. The evolution of the output error $\mathcal{E}$ then satisfies,
$\frac{d\mathcal{E}}{dt}=-2\Theta\mathcal{E},\quad\Rightarrow\quad\mathcal{E}(t)=\mathcal{E}(\theta_{0})e^{-2\Theta t}.$ (3)
This resembles a first-order rate equation governing exponential decay, with $\Theta$ acting as the decay factor. To isolate the eigenmodes and their respective decay rates (eigenvalues), the NTK square matrix can be diagonalized as $\Theta=\mathbf{Q}^{\top}\Lambda\mathbf{Q}$. Substituting the diagonalized form into Eqn. 3 and simplifying yields, $\mathcal{E}(t)=\mathbf{Q}^{\top}e^{-2\Lambda t}\mathbf{Q}\mathcal{E}(\theta_{0})$. This equation implies that the components of reconstruction error $\mathcal{E}$ associated with larger eigenvalues decay more rapidly than those associated with smaller eigenvalues. Applying the Fourier transform to the error evolution equation yields:
$\hat{\mathcal{E}}(\theta_{t})=\mathcal{E}(\theta_{0})\hat{\mathbf{Q}}^{\top}e^{-2\Lambda t}\hat{\mathbf{Q}}$. (4)
where $\hat{\mathbf{Q}}$ denotes the Fourier-transformed eigenbasis of the NTK, and $\hat{\mathcal{E}}$ represents the error expressed in the frequency domain. This equation implies that, the Fourier components of error $\mathcal{E}$ are correlated to Fourier components of NTK eigenvectors $\hat{\mathbf{Q}}$.
Figure 3: Contrasting learning dynamics of SIREN for low and high frequency-dominant targets. (a) Output evolution when fitting the low-frequency signal of Eqn. 5. The right subplot shows the NTK spectral energy, $\log(1+\mathcal{S}(k))$, across frequency eigenmodes $k$ during training, exhibiting a steady increase that indicates effective learning. (b) Output and NTK spectral energy evolution for a high-frequency signal. The left subplot shows that SIREN fails to match the ground truth, while the right subplot reveals a suppression of spectral energy over training, indicating difficulty in representing high-frequency-dominant signals.
Based on this premise, next we consider a toy problem where we attempt to supervise SIREN to fit two signals, one with predominantly low-frequency content, and another signal with predominantly high-frequency content. The signals are defined as follows:
$f(t)=\begin{cases}\sum\limits_{i=1}^{3}A_{k}\sin(2\pi k_{i}^{(L)}t),&\text{(Low-frequency signal)}\\ \sum\limits_{i=1}^{3}A_{k}\sin(2\pi k_{i}^{(H)}t),&\text{(High-frequency signal)}\end{cases}$ (5)
The amplitudes are normalized as $A=[\frac{1}{7},\frac{2}{7},\frac{4}{7}]$, and the frequencies are defined by $k_{i=1,3}^{(L)}=[0.25\tilde{k},\;0.50\tilde{k},\;0.75\tilde{k}$], $k_{i=1,3}^{(H)}=[0.85\tilde{k},\;0.90\tilde{k},\;0.95\tilde{k}$], with Nyquist frequency $\tilde{k}=\frac{N}{2}$ and $N=2^{16}$ samples.
We use a five-layer SIREN with 128 hidden units, $\omega_{0}=30$, and inputs scaled by 10, giving a first-layer frequency of 300. The network is trained to fit the low- and high-frequency signals in Eqn. 5. During training, we track both the reconstructed signal and the cumulative eigenvalue-weighted NTK power spectrum:
$\mathcal{S}(k)=\sum_{i=1}^{n}\lambda_{i}\,\left|\widehat{\mathbf{v}}_{i}(k)\right|^{2},$ (6)
where $\lambda_{i}$ are NTK eigenvalues and $\widehat{\mathbf{v}}_{i}(k)$ are the Fourier coefficients of the corresponding eigenvectors. $\mathcal{S}(k)$ measures the contribution of each frequency mode to the network’s representation during training.
Figure 4: Poor frequency support of SIREN for fitting high-frequency dominant targets. Distributions (column-1) and cumulative power spectra (column-2) of hidden-layer pre-activation and network outputs for a four-layer SIREN at epoch 0 and $10^{4}$ when fitting tetris.wav. Across all layers, the spectral content is concentrated far below that of the high-frequency target, indicating insufficient frequency support.
As shown in Fig. 3, SIREN accurately reconstructs the low-frequency target, with $\mathcal{S}(k)$ progressively increasing in the relevant modes. However, for the high-frequency target, the output remains near zero and $\mathcal{S}(k)$ is steadily suppressed, indicating an inability to fit high-frequency components. To address this phenomenon, we introduce a weight perturbation scheme in section IV.
Frequency support. We characterize the distributions and frequency response of pre-activations via the cumulative power spectral density (PSD) $\mathrm{PSD}(k)=\sum_{j=1}^{N_{h}}|\widehat{x}_{\mathrm{pre},j}(k)|^{2}$ of hidden-layer pre-activations. This experiment included the fitting of high-frequency data of tetris.wav. The signal contains 150,000 uniformly spaced samples in [1,1][-1,1] and is evaluated using a four-layer SIREN. Fig. 4 shows that across all hidden layers, the value distributions remain centered and narrow, while the cumulative PSD is heavily biased toward low frequencies. The network outputs, along with the PSD of intermediate-layer pre-activations, exhibit a spectral profile that deviates significantly from the ground-truth spectrum, which is dominant of high-frequencies. Even after $10^{4}$ epochs, the mismatch persists, revealing a fundamental limitation: the effective frequency support of SIREN pre-activations falls far off from the target spectrum. This restriction acts as a learning bottleneck caused by the mismatch of spectral profiles of the network pre-activations and the target. We term such a scenario as ‘spectral bottleneck’ and prevents the network from reconstructing signals like tetris.wav.
Remark. As observed in Figs. 3 and 4, an extreme case of a spectral bottleneck is characterized by: (a) progressive attenuation of the NTK eigenmode energy with each training epoch, (b) failure of the network to represent even the low-frequency components (highlighted in Fig. 4) that lie within the nominal frequency support of SIREN, and (c) convergence to a zero-valued output.
IV WINNER: Weight Initialization with Noise for NEural Representations
A weight perturbation scheme is proposed to address the spectral bottleneck of SIREN observed under its default weight initialization scheme sitzmann2020implicit. Traditional weight initializations, such as He and Glorot, aim to maintain the variance of pre- and post-activations within controlled bounds to avoid exploding or vanishing gradients, typically by scaling the weights inversely with the fan_in or fan_out. For example, the Glorot-style initialization used by Sitzmann et al. sitzmann2020implicit samples weights as $W_{jk}\sim\mathcal{U}\left(-\frac{1}{\omega_{0}}\sqrt{\frac{6}{\text{fan\_in}}},\frac{1}{\omega_{0}}\sqrt{\frac{6}{\text{fan\_in}}}\right)$ to ensure unit standard deviation for all pre-activations ($\mathbf{W}\cdot\mathbf{x}+\mathbf{b}$) of the SIREN.
In Sec. III-B, we have emphasized that the root cause of the spectral bottleneck (Fig. 3) is the mismatch of spectral energy between the target signal and network activations at initialization. We address this issue using Proposition 1 to alter the Fourier representation of the network’s activations and outputs at initialization. To this end, we introduce WINNER, a weight perturbation scheme in which a Gaussian noise is added to the uniformly initialized weights that exist between the inputs and the second hidden layer. So, the weights immediately upstream of first and second hidden layers are perturbed as,
$W_{jk}^{(l)}\leftarrow W_{jk}^{(l)}+\eta_{jk}^{(l)},$ (7)
where the noise matrix $\eta_{jk}^{l}$ is sampled from a normal distribution,
[…]
The Gaussian scale parameters $s_{0},s_{1}]$ control the width of the pre-activation distributions and their spectra. Based on the WINNER perturbation scheme, we introduce SIREN2, a perturbed variant of SIREN, (the extra N in N2 denotes Noise).
$\begin{gathered}\text{SIREN}^{2}:\;f(\mathbf{x};\theta)=\mathbf{W}^{(L)}\mathbf{h}^{(L-1)}+\mathbf{b}^{(L)},\\[4.0pt] \mathbf{h}^{(l)}=\begin{cases}\mathbf{x},&l=0\text{ (inputs)},\\[3.0pt] \phi^{\text{sin}}\!\big((\mathbf{W}^{(l)}+\eta_{jk}^{(l)})\mathbf{h}^{(l-1)}+\mathbf{b}^{(l)}\big),&l=1,2,\\[3.0pt] \phi^{\text{sin}}\!\big(\mathbf{W}^{(l)}\mathbf{h}^{(l-1)}+\mathbf{b}^{(l)}\big),&l=3,\dots,L-1.\end{cases}\end{gathered}$ (9)
Although the weight perturbations are added only up to the second hidden layer, their effect propagates downstream all the way to the outputs. This is shown empirically in Sec. V and in Supplementary Material Sec. LABEL:suppl:fulldists for the full network. The goal of the proposed noise addition scheme is to enhance the functional sensitivity between the outputs and network parameters necessary to allow the parameter updates required for fitting high-frequency modes.
Figure 5: Distributions of the dot product between inputs $X_{ij}\sim\text{Arcsin}(-1,1)$ and weights initialized under two initialization schemes: (a) standard uniform weights $W_{jk}=\sim\mathcal{U}\left(-\omega_{0}^{-1}\sqrt{6/d},\omega_{0}^{-1}\sqrt{6/d}\right)$, and (b) WINNER, in which uniform weights are perturbed with Gaussian noise, $W_{jk}+\eta_{jk}$. The noise addition increases the standard deviation of the dot product from $1$ to $\sqrt{1+\frac{ds^{2}}{2}}$, where $d$ is the input dimension (fan_in). This increase closely matches the analytically predicted value (Theorem IV.1) shown by the dashed black line.
[…]
Theorem IV.1.
Let the following matrices be defined:
  • Input Matrix $X\in\mathbb{R}^{n\times d}$: Each entry $X_{ij}$ is independently sampled from an arcsine distribution $\mathcal{A}(-1,1)$, which has a mean of 0 and a variance of $1/2$.
  • Weight Matrix $W\in\mathbb{R}^{d\times d}$: Each entry $W_{jk}$ is independently sampled from a uniform distribution $\mathcal{U}\left(-\frac{\sqrt{6/d}}{\omega_{0}},\frac{\sqrt{6/d}}{\omega_{0}}\right)$ for a given $\omega_{0}>0$.
  • Noise Matrix $\eta\in\mathbb{R}^{d\times d}$: Each entry $\eta_{jk}$ is independently sampled from a normal distribution $\mathcal{N}(0,(s/\omega_{0})^{2})$ for some scale parameter $s>0$.
Consider the perturbed matrix $\omega_{0}\cdot X(W+\eta)=Y^{\prime}$, where $Y^{\prime}\in\mathbb{R}^{n\times d}$. Then, for each entry $Y^{\prime}_{ik}$ of the matrix $Y^{\prime}$, its distribution is approximately Gaussian with zero mean and standard deviation $\sqrt{1+\frac{ds^{2}}{2}}$.
Proof.
The random distribution $Y^{\prime}$ can be decomposed as $Y^{\prime}=Y_{1}+Y_{2}$, with $Y_{1}=\omega_{0}(XW_{jk})$ and $Y_{2}=\omega_{0}(X\eta)$. Since $X$ and $W_{jk}$ are independent, the entries of $Y_{1}$ are sums of products of independent zero-mean random variables $X_{ij}$ and $W_{jk}$. Following sitzmann2020implicit (see their Theorem 1.8) and by the central limit theorem (CLT), each $Y_{1,ik}$ is approximately Gaussian with mean and variance $\mathbb{E}[Y_{1,ik}]=0$ and $\mathrm{Var}[Y_{1,ik}]=1$ respectively. Similarly, for $Y_{2,ik}=\omega_{0}\sum_{j=1}^{d}X_{ij}\eta_{jk}$, since $X_{ij}$ and $\eta_{jk}$ are independent and have zero-mean,
$\mathrm{Var}[X_{ij}\eta_{jk}]$ $=\mathbb{E}[(X_{ij}\eta_{jk})^{2}]-(\mathbb{E}[X_{ij}\eta_{jk}])^{2}$
$=\mathbb{E}[X_{ij}^{2}]\cdot\mathbb{E}[\eta_{jk}^{2}]-0=\frac{1}{2}\cdot\left(\frac{s^{2}}{\omega_{0}^{2}}\right).$
Now, for the sum,
$\sum_{j=1}^{d}\mathrm{Var}\left[X_{ij}\eta_{jk}\right]=\sum_{j=1}^{d}\frac{s^{2}}{2\omega_{0}^{2}}=d\cdot\frac{s^{2}}{2\omega_{0}^{2}}.$
Since $Y_{2,ik}=\omega_{0}X_{ij}\eta_{jk}$, the variance scales by a factor of $\omega_{0}^{2}$, yielding,
$\mathrm{Var}[Y_{2,ik}]=\omega_{0}^{2}\cdot\mathrm{Var}[X_{ij}\eta_{jk}]=\frac{ds^{2}}{2}.$
Finally, since $Y_{1}$ and $Y_{2}$ are independent, their variances add:
$\mathrm{Var}[Y^{\prime}_{ik}]=\mathrm{Var}[Y_{1,ik}]+\mathrm{Var}[Y_{2,ik}]=1+\frac{ds^{2}}{2}.$
$\Rightarrow\mathbb{E}[Y^{\prime}_{ik}]=0,\quad\mathrm{Std}[Y^{\prime}_{ik}]=\sqrt{1+\frac{ds^{2}}{2}}.$
Proposition 2.
Given that the standard deviation of pre-activations in the layer-2 of a SIREN scale by a factor of $\sqrt{1+\frac{ds^{2}}{2}}$ under the weight perturbation scheme in Eqn. 7, adding white noise $\eta_{jk}$ to the uniform weights $W_{jk}\sim\mathcal{U}\left(-\frac{1}{\omega_{0}}\sqrt{\frac{6}{\text{fan\_in}}},\frac{1}{\omega_{0}}\sqrt{\frac{6}{\text{fan\_in}}}\right)$ is approximately equivalent to scaling the activation frequency $\omega_{0}$ by the same factor. This approximation holds under the assumption that the contribution of the bias vector to the pre-activation statistics is negligible, which is justified for large $d$ (fan_in) values.
IV-A Target-aware Specification of Noise Scales $s_{0}$ and $s_{1}$
As shown in Figs. 1, 2, and 3, the performance of INRs is sensitive to the spectral content of the target signal. Therefore, an effective weight initialization is one that is target-aware and accounts for the spectral profile of the target. To quantify the weighted contribution of different frequencies to the target, a spectral centroid $\psi$ is defined as the normalized average frequency of the target’s power spectrum, computed as
$\psi=2\times\frac{\sum_{k}k|\hat{y}(k)|}{\sum_{k}|\hat{y}(k)|},$ (10)
where $\hat{y}(f_{k})$ is the Fourier transform of the target evaluated at frequency bin $k$. The factor $2$ normalizes $\psi$ to the range $[0,1]$. Using $\psi$, the noise scales $s_{0}$ and $s_{1}$ in Eqn. 8 are empirically determined as
$s_{0}=s_{0}^{\text{max}}\!\left(1-e^{-a\frac{\psi}{C}}\right),\quad s_{1}=s_{1}^{\text{max}}\left(\frac{\psi}{C}\right),$ (11)
where $C$ denotes the number of channels (e.g., $C=3$ for RGB images). The hyperparameters $s_{0}^{\text{max}},a,s_{1}^{\text{max}}]$ are set as $\left[\frac{3500}{70^{N_{d}-1}},5,\frac{4}{10^{N_{d}-1}}\right]$, with $N_{d}$ representing the input dimension of the network. Eqn. 11 is valid for all modalities including audio, images, video, and 3D data.
The functional form of $s_{0}$ in Eqn. 11 is empirically determined through multiple experiments. An initial linear formulation was tested, but experiments relating the spectral centroid to optimal $s_{0}$ values revealed a trend better captured by an exponential curve rather than a linear one (see Supplementary Fig. S1). The linear dependence was retained for $s_{1}$. Both expressions are therefore ad-hoc and derived from multiple empirical evaluations; further work is required to develop a principled, target-aware formulation. Additional plots illustrating the sensitivity of noise scales to fitting accuracy for various audio clips are included in the supplementary material. As also detailed in sitzmann2020implicit, the larger input scaling factor or the larger noise scales used for audio signals arise from their high sampling rates and higher frequency content compared to visual data.
Fig. 6 reports the effect of varying $s_{0}$ and $s_{1}$ on audio and image reconstruction. We find that SIREN2 maintains strong performance over a broad range of values, showing that the method is not overly sensitive to precise tuning. The cross-marked settings from Eqn. 11 reliably fall in regions of high PSNR, providing a simple and robust rule for setting the perturbation scales without expensive hyperparameter search.
Figure 6: Sensitivity to noise scales. PSNR for audio (left) and image (right) reconstruction as a function of the perturbation scales $s_{0}$ and $s_{1}$. Performance is stable across a wide range, with the cross-mark indicating the target-aware scales from Eqn. 11 that consistently yield near-optimal results.
SIREN2 vs. FINER++: The current work compares against FINER/FINER++ liu2024finer; zhu2024finerplusplus, which also aim to improve high-frequency representation but do so by modifying the bias initialization and the activation function. However, altering biases does not mathematically guarantee Gaussian-distributed pre-activations, and therefore cannot inherit the properties of Gaussian initialization that stabilize optimization and control signal propagation glorot2010understanding. Another key difference is that FINER++ does not modify the weights connecting the first and second hidden layers, which limits its ability to adjust the initial spectral profile of pre-activations downstream of the first hidden-layer.
V Spectral Properties of SIREN2
V-A Neural Tangent Kernel at Initialization
The spectral characteristics of the NTK at initialization are examined with an aim to understand the frequency support of SIREN2 in comparison to SIREN. As shown in Fig. 7, SIREN exhibits rapid eigenvalue decay and allocates higher cumulative spectral energy $\mathcal{S}(k)$ to the low-frequencies. This rapid collapse of energy occurs due to the alignment of dominant NTK eigenvectors with smooth, low-frequency functions, restricting expressivity in tasks requiring fine-scale resolution such as audio representation. The proposed WINNER used in SIREN2 provides significant improvement, featuring a tunable $\mathcal{S}(k)$ profile controlled by weight perturbation scales $s_{0}$ and $s_{1}$, analogous to the Fourier scales used in random Fourier embeddings rahimi2007random; tancik2020fourier. By contrast, SIREN2 captures high-frequencies and positional information directly through its sinusoidal activations, avoiding the quadratic increase in parameters required by random Fourier or positional embeddings to represent similar frequency content. As demonstrated in Fig. 7, SIREN2 can be configured to exhibit slower eigenvalue decay with higher cumulative spectral energy $\mathcal{S}(k)$ distributed across the entire spectrum with parameters $s_{0}$ and $s_{1}$. The critical factor for a successful implicit representation is the appropriate selection of $s_{0}$ and $s_{1}$ (Sec. IV-A).
Figure 7: Controlling the NTK spectra via noise scales $s_{0}$ and $s_{1}$ of WINNER. (a) Normalized NTK kernels for SIREN and SIREN2 with $s_{0},s_{1}]=[10,0.5]$, showing reduced off-diagonal correlations. (b) NTK eigenvalue decay profile (left) and eigenvalue-weighted FFT magnitude spectra - $\mathcal{S}(k)$ (right) for varying noise scales, demonstrating that SIREN2 broadens frequency support. All networks shown here employ four hidden layers with 256 features, use $\omega_{0}=30$, and are evaluated on $2^{10}$ uniformly sampled inputs in [1,1][-1,1].
V-B Activation Spectra
Figure 8 compares the distributions and power spectral densities (PSDs) of network inputs and the pre- and post-activation values in the first two hidden layers for SIREN and SIREN2. For SIREN, pre-activation distributions are approximately Gaussian, while post-activations follow the $\text{Arcsin}(-1,1)$ law, consistent across layers. SIREN2 preserves these distributional structure, albeit with broader pre-activation spreads in layers 1 and 2. Distributions for subsequent layers (not shown) remain identical to that of SIREN; full-layer distributions are reported in the Supplementary (Sec. LABEL:suppl:fulldists).
Although similar distributions in the real space, differences arise in the spectral domain, especially the shape of spectra, due to the noise scales $s_{0}$ and $s_{1}$. SIREN exhibits dominant low-frequency content with negligible excitation of higher modes across layers, consistent with spectral bias. In contrast, SIREN2 introduces structured broadband spectral energy right from the first hidden layer, which propagate through subsequent layers up to the output. This leads to a sustained high-frequency energy in the activations, enhancing the conditioning of the optimization landscape for regressing high-frequency signals. SIREN2 retains the favorable bell type distributional properties of SIREN while enabling superior high-frequency receptivity at initialization.
Figure 8: WINNER enhances high-frequency receptivity. Input, pre-activation, and post-activation distributions are shown alongside layer-averaged power spectral densities (PSDs) of SIREN and SIREN2 up to layer-2 at initialization. SIREN2 exhibits higher variance in its pre-activation distributions due to weight perturbations introduced by WINNER ($s_{0}=100,s_{1}=1$). The PSDs reveal that SIREN has limited high-frequency content, whereas SIREN2 maintains broader spectral coverage with larger amplitudes at higher frequencies. This behavior persists through depth and extends to the outputs. A detailed analysis for a four-layer network is provided in the supplementary material. Both models use four hidden layers with 2048 hidden units per layer and are evaluated on $2^{10}$ uniformly spaced inputs over $\mathbf{x}\sim\mathcal{U}(-1,1)$ with $\omega_{0}=30$.
VI Experiments
We evaluate the performance of SIREN2 (SIREN initialized with WINNER) against several state-of-the-art INR architectures from the literature, including the baseline SIREN sitzmann2020implicit, Gauss ramasinghe2022beyond (2022), WIRE saragadam2023wire (2023), and FINER liu2024finer (2024), in reconstructing a variety of challenging audio signals. Following the recommendation in sitzmann2020implicit, all architectures, except SIREN2, use a scaling factor of 100 for the first-layer activation periodicity for the audio fitting experiments.
VI-A 1D Audio Fitting
To ensure consistency across experiments, all audio signals are fixed at a length of 150,000 samples, making the experimental setup invariant to sampling rate. Each network is trained for 30,000 epochs, and every experiment is conducted over five random trials to compute the mean peak PSNR and standard deviation. The selected audio signals span a range of spectral characteristics, including high-frequency-dominant, low-frequency-dominant, and broadband signals.
Table I shows that the proposed SIREN2 delivers consistently higher reconstruction accuracy than existing INR architectures across diverse audio signals, establishing new state-of-the-art results. These gains are especially notable for signals with strong high-frequency content, where other methods exhibit substantial residual errors. While the weight perturbation scheme amplifies high-frequency energy (Fig. 5), it slightly suppresses low frequencies, as seen in the reconstruction of bach.wav. This indicates the need for a more robust choice of $s_{0}$ and $s_{1}$ for such low-frequency dominant signals. As illustrated in Fig. 9, SIREN2 enables high-fidelity audio reconstructions, achieving PSNR values above $60$ dB with parameter count comparable to that of number of samples in the signal.
TABLE I: Audio fitting. Mean and standard deviation of PSNR values for different architectures on audio signal reconstruction. The proposed SIREN2 achieves state-of-the-art performance. Results are color coded as best, second best, and third best reconstructions. The network width is chosen so that the total parameter count is approximately equal to the signal length.
SIREN FINER WIRE ReLU-PE SIREN-RFF FINER++ SIREN2 (present)
[…]
# Fourier features 0 0 0 193 193 0 0
# parameters 149185 149185 149474 150155 150155 149185 149185
PSNR (dB) ($\uparrow$):
[…]
Figure 9: High-fidelity audio reconstruction with SIREN2. Each column corresponds to an audio signal: tetris.wav, sparking.wav, and shattered.wav. Top row: ground truth spectrograms. Second row: Averaged PSNR histories from different network architectures with the region of standard deviation shaded. Remaining rows show the spectrogram errors for reconstructions by each model. SIREN2 stands out by exhibiting near-zero reconstruction noise.
Reproducibility details. The inputs and audio targets are normalized to [1,1][-1,1] before training. SIREN and FINER use $\omega_{0}=30$, WIRE uses $\omega_{0}=10$ and $s_{0}=10$ saragadam2023wire, and SIREN-RFF uses $\omega_{0}=30$ with Fourier embeddings drawn from $\mathcal{N}(0,30^{2})$. FINER++ zhu2024finerplusplus employs a first-layer bias uniformly distributed in [5,5][-5,5]. For all networks, the first-layer $\omega_{0}$ is scaled by $100$. Training uses a learning-rate scheduler that decays by 1% every 20 epochs from an initial value of $10^{-4}$. Audio samples and code are provided in the linked GitHub repository.
VI-B 2D Image Fitting
We evaluate SIREN2 on 2D image fitting tasks, $f(\mathbf{x};\theta):\mathbb{R}^{2}\mapsto\mathbb{R}^{d}$, with $d=1$ for grayscale and $d=3$ for RGB images. The experiments cover a diverse set of images, including natural images from the Kodak kodak_photocd dataset, challenging texture images from two DTD cimpoi14describing classes (braided and woven), and synthetic high-frequency patterns. Table II shows that SIREN2 consistently outperforms the original SIREN across all cases, with PSNR improvements ranging from 7% (2D-Riemann) to 69% (noise.png), and especially strong gains for images with high-frequency content and small pixel count (in the over-paramaterized regime). For RGB image reconstruction, SIREN2 provides only marginal gains over SIREN, while FINER achieves the best performance. Figure 10 visualizes these improvements using FFT error maps, highlighting that SIREN2 achieves lower fitting error in high-frequency regions. The frequency-domain analysis further confirms that SIREN2 preserves fine details and sharp transitions more accurately, as indicated by the reduced magnitude errors (darker regions).
TABLE II: 2D image fitting. PSNR($\uparrow$) in dB across images and datasets for different INR architectures. SIREN2 consistently surpasses SIREN, with percentage gains (in parentheses) reflecting improvements achieved purely through initialization. Results are color coded as best, second best, and third best reconstructions.
SIREN SIREN2 (present) ReLU-PE WIRE FINER Gauss
[…]
# parameters 198145 198145 263553 198386 198145 198145
Peak PSNR (dB) ($\uparrow$):
noise.png 21.3 36.1 (69% $\uparrow$) 16.9 25.5 33.0 34.1
camera.png 38.9 44.9 (15% $\uparrow$) 28.4 37.2 46.4 28.6
castle.jpg 33.6 36.5 (9% $\uparrow$) 22.3 28.5 36.9 19.2
rock.png 26.9 36.2 (35% $\uparrow$) 16.1 26.5 36.6 31.8
2D-Riemann (CFD data) 55.3 59.1 (7% $\uparrow$) 45.8 49.7 60.5 28.1
DTD braided dataset (120 images, gray mode) 48.6 75.2 (55% $\uparrow$) - - 65.4 -
DTD woven dataset (120 images, gray mode) 41.9 61.0 (46% $\uparrow$) - - 53.1 -
Kodak dataset (24 images, gray mode) 34.9 37.6 (8% $\uparrow$) - - 38.1 -
The reported PSNR values for these datasets represent averages computed over all images.
Figure 10: Accuracy improvements in Fourier space with SIREN2 for image fitting. Reconstruction performance of SIREN and SIREN2 models for various images. The top two rows show the ground truth images and their corresponding FFT magnitudes ($|\mathcal{F}(f_{gt})|$). Subsequent rows depict the FFT error maps and peak PSNR values achieved by different models. SIREN2 consistently produce lower FFT errors and higher PSNRs.
Reproducibility details. All networks follow the same settings as the audio fitting experiments, with no scaling applied to the first layer $\omega_{0}$. The ReLU+PE presented in Table II incorporates positional encoding with 256 embeddings. The positional encodings use a logarithmic frequency spectrum, with frequencies ranging from $2^{0}$ to $2^{\text{n}-1}$, with n=7 (number of frequencies).
VI-C Image Denoising
Figure 11: Image denoising. Comparison of reconstructions for noisy 2D fields and natural images. Rows 1–2: reconstruction of a noisy density field obtained from a computational fluid dynamics simulation of a 2D Riemann problem kurganov2002solution; chandravamsi2024high. Rows 3–4: reconstruction of a sparrow image and a nebula image. SIREN2 yields the highest fidelity among the compared methods, with quantitative results summarized in Table III.
The robustness of different INR architectures is assessed for the canonical image denoising. Firstly, a clean signal $f(\mathbf{x})$ is corrupted by additive white Gaussian noise $\eta(\mathbf{x})\sim\mathcal{N}(0,\sigma^{2})$ such that $\tilde{f}(\mathbf{x})=f(\mathbf{x})+\eta(\mathbf{x})$ achieves a signal-to-noise ratio of $\mathrm{SNR}=5\ \mathrm{dB}$. The task is to reconstruct $f$ from $\tilde{f}$. We adopt an unsupervised denoising strategy similar to Noise2Self batson2019noise2self, training the INR directly on the noisy input while reserving a small subset of pixels for ‘J-invariant’ validation. A predictor is J-invariant if, for any pixel $i$, the prediction at $i$ does not depend on the noisy value at $i$ itself. This restriction forces the model to reconstruct structure from spatial context rather than memorize pixelwise noise. The held-out set then provides an unsupervised early-stopping criterion that prevents overfitting to noise while retaining underlying structure. Ground-truth images are used strictly for evaluation (e.g., PSNR, SSIM) and never for model selection. For the image quality evaluation, along with PSNR we report structural similarity wang2004image, mean absolute error (MAE), LPIPS zhang2018unreasonable, and DISTS ding2020image. For clarity, arrows ($\uparrow$ / $\downarrow$) are used in tables and figures to indicate whether higher or lower values correspond to better performance.
TABLE III: Quantitative comparison of different architectures on noisy image reconstruction. SIREN2 (ours) consistently achieves the best or comparable scores across all metrics, showing better noise suppression and structural fidelity.
PSNR($\uparrow$) SSIM($\uparrow$) MAE($\downarrow$) LPIPS($\downarrow$) DISTS($\downarrow$)
2D Riemann (Fig. 11)
Ground truth + Noise 20.48 0.133 0.0734 0.6836 0.3078
SIREN 32.62 0.802 0.0173 0.4139 0.2732
FINER 32.58 0.790 0.0176 0.4397 0.2686
WIRE 28.06 0.449 0.0314 0.6005 0.2864
SIREN2 (Ours) 32.98 0.812 0.0164 0.4096 0.2648
Sparrow
Ground truth + Noise 20.44 0.144 0.0753 0.7077 0.2722
SIREN 29.49 0.905 0.0181 0.4417 0.1816
FINER 31.90 0.880 0.0172 0.4929 0.1423
WIRE 30.23 0.816 0.0205 0.5336 0.1792
SIREN2 (Ours) 32.14 0.883 0.0167 0.4899 0.1414
Nebula
Ground truth + Noise 20.31 0.261 0.0775 0.5554 0.2196
SIREN 25.04 0.664 0.0297 0.5747 0.2419
FINER 26.37 0.697 0.0263 0.5055 0.1503
WIRE 25.02 0.582 0.0350 0.5410 0.1997
SIREN2 (Ours) 26.97 0.695 0.0259 0.4827 0.1120
Table III summarizes the quantitative evaluation of image denoising performance across different models using standard metrics, including PSNR, SSIM, MAE, LPIPS, and DISTS. The results demonstrate that SIREN2 consistently achieves higher fidelity reconstructions with improved perceptual and structural quality compared to the baseline SIREN and other methods. Snapshots of reconstructions presented in Fig. 11 with regard to the 2D-Riemann problem suggest that SIREN and FINER oversmooth fine-scale details, while WIRE preserves global appearance but leaves residual noise. SIREN2 achieves the sharpest and most faithful reconstructions (with the best SSIM and LPIPS), suppressing noise while retaining textures and edges.
Reproducibility details. The ground truth signal is corrupted with Gaussian noise at an SNR of 5 dB. All experiments employ a four-layer architecture with 256 features per layer. The search space for SIREN2 noise scales is confined to $s_{0}\in[0,200]$ with a fixed $s_{1}=0.01$. For FINER++, the bias scale $k$ in $\bar{b}\sim\mathcal{U}(-k,k)$ of the first hidden layer is set to $k\in[0,20]$. The first-layer activation periodicity $\omega_{0}$ is kept at 30 in both FINER++ and SIREN2, ensuring comparable spectral bias.
VI-D Audio Denoising
TABLE IV: Audio denoising. Best PSNR ($\uparrow$) in dB for different audio clips using FINER++ and SIREN2.
GT+Noise FINER++ SIREN2 (present)
bach.wav 21.16 34.69 35.53
dilse.wav 21.46 35.16 35.18
birds.wav 29.92 34.84 37.17
counting.wav 26.67 38.19 38.71
We adopt the same Noise2Self-inspired DIP training procedure from Sec. VI-C for the present audio denoising experiments. Ground-truth audio signal is used only for evaluation. While the general denoising framework is identical, audio signals present distinct challenges. Unlike natural images, which concentrate most energy in low frequencies, audio signals often exhibit a relatively broadband structure (e.g., higher harmonics in music, ambient noise), causing stronger overlap between the underlying signal and Gaussian noise. This overlap makes simple frequency-selective filtering less effective. In an a-priori setting, where no ground-truth information is available, INR-based denoising can provide a useful alternative.
in Sec. VI-A on supervised audio fitting task, $\omega_{0}$ was scaled by a factor of 100 consistently for all network architectures. However, for the present audio-denoising task, a large first-layer periodicity consistently degraded denoising for SIREN, Gauss, WIRE, and FINER, with best denoising $\mathrm{PSNR}<25\,\mathrm{dB}$ for different signals; due to the presence of broadband noise in the noisy signal. We therefore avoid first-layer $\omega_{0}$ scaling for denoising and report results for FINER++ and SIREN2 architectures only which does not use any first layer $\omega_{0}$ scaling. Table IV and Fig. 12 report the accuracy of denoised reconstructions for bach.wav audio clip after 20,000 epochs for FINER++ and SIREN2. The noise scales $s_{0}$ and $s_{1}$ in SIREN2 provide controllable filtering, allowing a principled adjustment of frequency support to match the noise characteristics. For FINER++, varying the bias ranges was evaluated. The results in Table IV show that both networks achieve competitive accuracy in the recovery of the underlying signal.
Reproducibility details. The ground truth signal is corrupted with Gaussian noise at an SNR of 5 dB. All experiments employ a four-layer architecture with 256 features per layer. The search space for SIREN2 noise scales is confined to $s_{0}\in[800,2000]$ with a fixed $s_{1}=0.001$. For FINER++, the bias scale $k$ in $\bar{b}\sim\mathcal{U}(-k,k)$ of the first hidden layer is set to $k\in[0,20]$. The first-layer activation periodicity $\omega_{0}$ is kept at 30 in both FINER++ and SIREN2. No scaling factor was used to increase the $\omega_{0}$ value for the first layer activations.
Figure 12: Audio denoising. Log-spectrogram comparison for audio denoising task performed on bach.wav using FINER++ and SIREN2. SIREN2 achieves higher PSNR and preserves more harmonic detail, as visible in the spectrogram. The rightmost plot shows the PSNR variation of SIREN2 over the $(s_{0},s_{1})$ grid, illustrating that the noise scales $s_{0}$ and $s_{1}$ provide flexibility to control the degree of denoising through modulation of high-frequency suppression.
VI-E Audio Inpainting
We evaluate the inpainting capability of SIREN2 using the same network width, signal length, and noise scales employed in the audio fitting experiments of Sec. VI-A. A fixed percentage of sample values is removed and the model is trained only on the observed subset. Sparsity denotes the retained fraction of samples in the training set. No spatial or structural priors are used; the task reduces to regressing the underlying continuous audio data from incomplete pointwise samples.
Across all sparsity levels, SIREN2 converges to lower reconstruction error for most cases than the baseline SIREN. The gains are consistent for both low and moderate sparsity, indicating that the improved high-frequency receptivity at initialization can benefit recovery of missing regions. The reconstructed samples exhibit sharper transitions indicating influence of initialization scheme on the INR’s ability to approximate the fine-scale details under limited observations.
TABLE V: Audio inpainting. Mean and standard deviation of PSNR (in dB) for different sparsity levels using SIREN and SIREN2.
Audio sample Sparsity SIREN SIREN2
bach.wav 0.05 18.87 $\pm$ 0.28 19.19 $\pm$ 0.03
0.1 22.26 $\pm$ 0.14 22.71 $\pm$ 0.13
0.25 32.79 $\pm$ 0.63 32.97 $\pm$ 0.37
0.5 46.89 $\pm$ 0.33 46.81 $\pm$ 0.44
relay.wav 0.05 16.53 $\pm$ 1.51 18.34 $\pm$ 0.08
0.1 15.92 $\pm$ 0.28 18.10 $\pm$ 0.17
0.25 16.79 $\pm$ 0.07 18.18 $\pm$ 0.18
0.5 17.77 $\pm$ 0.03 19.18 $\pm$ 0.16
tetris.wav 0.05 8.80 $\pm$ 0.13 12.19 $\pm$ 0.02
0.1 9.21 $\pm$ 0.07 11.95 $\pm$ 0.03
0.25 10.13 $\pm$ 0.21 11.87 $\pm$ 0.08
0.5 12.19 $\pm$ 0.27 12.50 $\pm$ 0.08
VII Conclusion
Deep neural networks exhibit a spectral learning bias, where low-frequency components are learned early in training while higher-frequency modes emerge more gradually in the later epochs. However, when the target signal lacks low-frequency content and is dominated by high frequencies, training can suffer from a spectral bottleneck, leading to failure in reconstructing the signal. The present work examined such a phenomenon in the context of sinusoidal representation networks (SIRENs), with a focus on fitting signals dominated by high-frequency content. We analyze the evolution of activation spectra and the neural tangent kernel and characterize the mechanisms underlying the spectral bottleneck. To address this limitation, we propose a target-aware Gaussian weight perturbation scheme, WINNER. The method applies Gaussian noise of a specified scale to the uniformly initialized weights, thereby modifying pre-activation distributions and layerwise spectra. Similar to random Fourier embeddings tancik2020fourier, the present weight perturbation scheme can be used to control the empirical NTK and its eigenbasis, with the added benefit of not introducing additional trainable parameters. The approach achieves state-of-the-art performance on audio fitting tasks and highlights the critical role of initial activation spectra in determining fitting accuracy.
Limitations. (a) The proposed initialization scheme relies on prior knowledge or an approximate estimate of the target’s spectral content to determine the perturbation scales $s_{0}$ and $s_{1}$. For problems where the target is not known in advance, such as denoising or solving initial/boundary value PDEs, these hyperparameters can be initialized with estimated values and adaptively updated during early training. A similar limitation also arises when specifying the Fourier scale for other equivalent methods such as random Fourier feature embeddings tancik2020fourier. (b) Similar to the other architectures examined in this study, the training dynamics and final performance of SIREN2 are sensitive to the learning rate and its decay schedule. Without a scheduler, training often results in a strongly oscillatory PSNR evolution resulting in poor convergence. We recommend a decay rate of 1-2% every 20 epochs with an initial learning rate of $10^{-4}$ for both audio and image fitting to achieve stable, non-oscillatory PSNR evolution.
Acknowledgments
The authors HC, DVS, IZ, ZC, and SP gratefully acknowledge the financial assistance provided by Technion - Israel Institute of Technology during the course of present work. HC thanks Sanketh Vedula (Princeton University) for directing him to the work of Sitzmann et al. sitzmann2020implicit; experimenting with it helped the development of noise scheme presented in this work. Generative AI tools were used solely for language editing, limited to grammar and phrasing refinement, and were not used for developing or modifying technical content.
Biographies
Hemanth Chandravamsi received the Ph.D. degree in mechanical engineering from the Technion – Israel Institute of Technology, Haifa, Israel. He is currently a Postdoctoral Researcher with the Faculty of Mechanical Engineering at the Technion. His research interests include scientific and physics aware machine learning, high speed flows, and numerical algorithms.
Dhanush V. Shenoy received the Ph.D. degree in Aerospace Engineering from ISAE-SUPAERO, Toulouse, France, where his doctoral research focused on aeroacoustics and low–Reynolds-number flows. He is currently pursuing post-doctoral research in deep reinforcement learning and its applications to computational fluid dynamics. His research interests include machine learning for flow modeling and control, quantum computing and quantum-inspired algorithms, lattice Boltzmann methods, high-performance computing, and GPU-accelerated numerical solvers.
Itay Zinn received the M.Sc. degree in mechanical engineering from the Ariel University, Israel, and he is currently pursuing his Ph.D. degree in computational fluid dynamics at Technion – Israel Institute of Technology, Israel. His research interests include machine learning, high speed flows, and computational fluid dynamics.
Ziv Chen received the B.Sc. degree from the Technion – Israel Institute of Technology, Haifa, Israel, where he is currently pursuing a direct Ph.D. degree with the Faculty of Electrical and Computer Engineering. He is also affiliated with the Helen Diller Quantum Center, Technion – Israel Institute of Technology, from which he received an excellence scholarship. His research interests include scientific and physics aware machine learning and quantum computing algorithms for physics simulations.
Shimon Pisnoy received the Ph.D. degree in mechanical engineering from the Technion – Israel Institute of Technology, Haifa, Israel, where he is currently a Ph.D. student with the Faculty of Mechanical Engineering. His research interests include machine learning, high speed flows, and computational fluid dynamics.
Steven H. Frankel is a Rosenblatt Chair Professor with the Faculty of Mechanical Engineering, Technion – Israel Institute of Technology, Haifa, Israel. He is also affiliated with the Helen Diller Quantum Center, Technion – Israel Institute of Technology. His research interests include scientific machine learning, computational fluid dynamics, combustion, and quantum computing for physics simulations.